A liquid X of density 3.36 g/cm3 is poured in a U-tube, which contains Hg. Another liquid Y is poured in left arm with height 8 cm, upper levels of X and Y are same. What is density of Y?a. 0.8 9/00 b. 1.2 g/cc c. 1.4 9/00 d. 1.6 9/00

Grade:

6 Answers

Arun
5 years ago
Dear Sarthak
 
Hx * \rhox * g = Hy * \large \rhoy * g + Hhg * \large \rhohg * g
10 * 3.36 = 8 * \large \rhoy + 2 * 13.6
\large \rhoy = 0.8 gm/cm³
 
Regards
Arun (askIITiams forum expert)
Saurav
5 years ago
actually the questio is incomplete!
but. i googled it and found the complete question
the height of  X should also be given
it is 10cm there
now balance the pressur below 10cm from the top in both limbs of the u tube
you will get the following equation
d= density
dx * g * H=  dy * g * Hy + dmercury * g * Hmercury
as we are equating pressure below 10 cm from the top 
8cm will be of y and 2 cm of mercury in left limb
therefore Hmercury = 2cm
also dmercury=13.6 gm/cm3
3.36 * 10 = dy * 8 + 13.6*2
dy=0.8 gm/cm3
I WONDER HOW ARUN SOLVED IT WITHOUT HX!!!!!!!!!!!!!
Arun
5 years ago
Dear Saurav
Β 
I also thought the same as you have thought.
I solved it on my experience. I know that some thong is missing but can't reject by saying it incomplete question.
User
5 years ago
Β 
Β height of Β X should also be given
it is 10cm there
now balance the pressur below 10cm from the top in both limbs of the u tube
you will get the following equation
d= density
dxΒ * g * HxΒ = Β dyΒ * g * HyΒ + dmercuryΒ * g * Hmercury
as we are equating pressure below 10 cm from the topΒ 
8cm will be of y and 2 cm of mercury in left limb
thereforeΒ HmercuryΒ = 2cm
also dmercury=13.6 gm/cm3
3.36 * 10 = dyΒ * 8 + 13.6*2
dy=0.8 gm/cm3
Hence option A
Rishi Sharma
3 years ago
Dear Student,
After reading the complete question and missing data from the comments.
Please find below the solution to your problem.

Density of liquid X-3.36g/ml
Density of liquid Y-dy
Density of mercury - 13.6
height of liquid X is 10 cm
height of liquid y is 8cm
Hy. Dy. g= hx. dx. g +HmDmg
Dy=60.8/8
=0.77=0.8

Thanks and Regards
Anuj Jain
2 years ago
namaste’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’’toΒ  all